# How to Make the Most of LLMs’ Grammatical Knowledge for Acceptability Judgments

### v1

@@ Abstract
The grammatical knowledge of language models (LMs) is often measured using a benchmark of linguistic minimal pairs, where LMs are presented with a pair of acceptable and unacceptable sentences and required to judge which is acceptable. The existing dominant approach, however, naively calculates and compares the probabilities of paired sentences using LMs. Additionally, large language models (LLMs) have yet to be thoroughly examined in this field. We thus investigate how to make the most of LLMs’ grammatical knowledge to comprehensively evaluate it. Through extensive experiments of nine judgment methods in English and Chinese, we demonstrate that a probability readout method, in-template LP, and a prompting-based method, Yes/No probability computing, achieve particularly high performance, surpassing the conventional approach. Our analysis reveals their different strengths, e.g., Yes/No probability computing is robust against token-length bias, suggesting that they harness different aspects of LLMs’ grammatical knowledge. Consequently, we recommend using diverse judgment methods to evaluate LLMs comprehensively.111 Our codes and templates will be publicly available upon acceptance.
@@
Yusuke Ide    Yuto Nishida    Miyu Oba    Yusuke Sakai Justin Vasselli    Hidetaka Kamigaito    Taro Watanabe Nara Institute of Science and Technology {ide.yusuke.ja6, nishida.yuto.nu8, oba.miyu.ol2, sakai.yusuke.sr9, vasselli.justin_ray.vk4, kamigaito.h, taro}@is.naist.jp
@@ 1 Introduction
Acceptability judgments have been widely used to measure grammatical knowledge of language models (LMs) Lau et al. (2017); Warstadt et al. (2019). Of two major categories of acceptability judgment benchmarks, we focus on the minimal-pair (MP) benchmark, where LMs are tested to see if they will prefer the more acceptable sentence from a pair of minimally different sentences. An example minimal pair extracted from Warstadt et al. (2020) is shown below.
@@
Here, sentence (a) is acceptable or grammatically correct, while (b) is not, as its underlined verb violates the subject-verb agreement. As such, MP benchmarks can evaluate any LMs including ase models and nstruct models, without fine-tuning for acceptability judgments.
@@
Meanwhile, recent scaling up of model sizes and training data for LMs has made it possible to solve a wide range of tasks by few-shot or zero-shot prompting, without task-specific finetuning Brown et al. (2020); Liu et al. (2021), popularizing the term large language models (LLMs). Incorporating learning techniques such as instruction-tuning Wei et al. (2022) and Direct Preference Optimization (DPO) Rafailov et al. (2023) has further improved the alignment of LLM outputs with human preferences and expectations. The LLMs trained by such techniques achieve good performance through prompting. In other words, LLMs show high performance when provided with guidance on what knowledge to elicit.
In this light, one can conceive various methods of obtaining acceptability judgments from LLMs, including prompting. However, no previous studies have thoroughly explored them; most of them naively input the given sentences to an (L)LM, calculate their probabilities, and deem the sentence with the higher probability of the pair to be acceptable for the (L)LM. Consequently, it is unclear what methods are effective in obtaining acceptability judgments using LLMs and what their strengths or weaknesses are.
We thus investigate how to make the most of LLMs for acceptability judgments, comparing (1) conventional sentence probability readout222 Readout refers to accessing an LLM’s output layer to compute probabilities of strings Kauf et al. (2024). methods, (2) novel probability readout methods in in-template settings, and (3) prompting-based methods. In in-template probability readout, we insert each target sentence into a template before inputting it to an LLM, instructing the LLM to focus on its grammaticality. As prompting-based methods, we investigate Yes/No probability computing (Yes/No prob comp) as well as simple prompting. In Yes/No prob comp, we compute the normalized probability of “Yes” versus “No”, inspired by UniEval Zhong et al. (2022), which is shown to be strong in evaluating natural language generation. We conduct rigorous experiments and analysis using six state-of-the-art LLMs and two MP benchmarks (one for English and one for Chinese) to demonstrate the following key findings.
@@
Figure 1: Conceptual illustration of in-template LP and Yes/No probability computing. Differences between paired sentences are underlined. Details are in Section 3.
  1. 1.
    An in-template probability readout method, in-template LP, and Yes/No prob comp (See Figure 1 for their conceptual illustration) show top performance, surpassing the conventional methods.
  2. 2.
    In-template LP and Yes/No prob comp have different strengths; for example, Yes/No prob comp is robust against token-length bias. This indicates that they harness different aspects of LLMs’ grammatical knowledge, helping comprehensive evaluation of LLMs.
  3. 3.
    Ensembling the two methods further improves the accuracy, revealing their complementary capabilities. The highest score by Mix-P3 with Qwen2 is 1.6 percentage points higher than humans on the English benchmark.
  4. 4.
    Even with the top two methods, all the LLMs have trouble making correct judgments where the unacceptable sentence can be obtained by shuffling the words in the acceptable one.
In conclusion, we recommend employing diverse judgment methods instead of relying on conventional sentence probability readout methods.
@@ 2 Related Work
2.1 Acceptability Judgments
Benchmarks of acceptability judgments have two categories, one of which is a single-sentence binary classification as seen in CoLA, a dataset composed of sentences each tagged as acceptable or unacceptable Warstadt et al. (2019). CoLA was incorporated into the natural language understanding benchmark GLUE Wang et al. (2018) and has been widely used to evaluate models. However, single-sentence benchmarks are limited in their ability to measure LMs’ grammatical knowledge directly because they require training a supervised classifier before the evaluation. This makes it difficult to distinguish between the knowledge of the model itself and what is learned through training the classifier Warstadt et al. (2020).
In contrast, MP benchmarks do not need task-specific training as they present minimally different pairs, asking which is acceptable. As another advantage of the MP benchmark, the minimal pairs are automatically generated in a controlled manner, which provides a sufficient amount of quality data for model evaluation Linzen et al. (2016). In conventional experiments using an MP benchmark, judgments are made based on sentence probabilities. Models are evaluated by whether they assign a higher probability to the acceptable sentence in each minimal pair. This method, which we call sentence probability readout, has been dominantly employed for MP acceptability judgments across languages (Marvin and Linzen, 2018; Warstadt et al., 2020; Mueller et al., 2020; Xiang et al., 2021; Someya and Oseki, 2023, inter alia).
As an exception in such studies, Hu and Levy (2023) compared the sentence probability readout and prompting. They conducted experiments using LLMs and an MP benchmark, which is composed of subsets of BLiMP Warstadt et al. (2020) and SyntaxGym Gauthier et al. (2020). They showed that sentence probability readout generally outperforms prompting. However, this study is limited in that their readout method remained conventional sentence probability readout and that instruction-tuned models were not investigated.
Another line of work has revealed that the token length influences the performance of sentence probability readout; normalized measures such as PenLP Wu et al. (2016) have been shown to mitigate some of this bias Lau et al. (2020), but they do not eliminate it Ueda et al. (2024).
@@ 2.1 Acceptability Judgments
Table 1: Example inputs of the readout methods. The target or inserted sentences are in italics.
@@
Table 2: Example messages for prompting. The target or inserted sentences are in italics.
@@ 3.1 Sentence Probability Readout
We experiment with three measures to compute the probability scores: LP, MeanLP, and PenLP. LP is the unnormalized log probability of the sentence
$\mathrm{LP}(\bm{s})$ $=\log P(\bm{s}).$ (1)
@@
where $\bm{s}$ is the input sequence of tokens and $P(\bm{s})$ is the probability assigned to $\bm{s}$ by the model. $\alpha$ is a hyperparameter to scale the token-length; we set $\alpha=0.8$ following Lau et al. (2020); Ueda et al. (2024). We hereafter refer to the three judgment methods simply by the name of the corresponding measures: LP, MeanLP, PenLP.
@@ 3.2 In-template Probability Readout
In-template probability readout follows the same steps of computing and comparing probabilities as sentence probability readout. Its input string, however, is built by embedding the sentences in a template designed to draw focus to their grammaticality. The input has two types: in-template single and in-template comparative. For each type, we prepare five templates per language because the performance can vary due to minor differences in expressions within prompts Zheng et al. (2023). The templates were created based on those of Flan333 https://github.com/google-research/FLAN Wei et al. (2022). For Chinese experiments, we use translation of English templates.
@@
The performance of in-template single inputs depends on the normalization, like the sentence probability readout. We thus apply each of the three measures explained above to the method, dubbing the corresponding methods in-template LP, in-template MeanLP, and in-template PenLP, respectively. The final measure also depends on whether we let $\bm{s}$ the whole input string or the target sentence only. We report the result of the former because it performed better in our preliminary experiments.
@@
In-template comparative inputs are built by filling two placeholders; we insert the target sentence into the first one and the other sentence of the minimal pair into the second. Note that the other sentence is supplementary, and the main aim here is to measure the acceptability of the target sentence.
Meanwhile, in-template comparative does not need normalization because the token length of the whole input string is constant no matter which of the paired sentences enter the first placeholder. We thus only calculate LP for the in-template comparative input, referring to this method as in-template comparative LP.
3.3 Prompting-based Methods
@@
A/B prompting inputs a prompt containing the paired sentences to the models and asks which sentence is acceptable. The prompt is exemplified in Table 2. The user message contains one acceptable and one unacceptable sentence. Their order (which sentence goes to A or B) is randomized to eliminate the potential bias from the order Pezeshkpour and Hruschka (2023). We perform constrained decoding by outlines555 https://github.com/outlines-dev/outlines Willard and Louf (2023) to ensure that the model outputs either A or B. 666 Our preliminary experiments without outlines observed many outputs violating the constraint. We turn off sampling in decoding.
@@
In Yes/No prob comp, we compute the score of each sentence as the normalized probability of “Yes” versus “No” given a prompt asking its acceptability. An example prompt is shown in Table 2. We predict the sentence that resulted in a higher “Yes” probability to be acceptable. This method is inspired by UniEval Zhong et al. (2022), which shows strong performance in evaluating natural language generation tasks. We formulate the probability given a sentence $\bm{s}$ as follows,
@@
where $P_{\mathrm{LLM}}(\cdot)$ is the probability of a token assigned by the model. For Chinese, we substitute “是” and “否” for “Yes” and “No”, respectively. In all our experiments, no tokenizers segment these words into subwords.
BLiMP CLiMP
Llama-3 Llama-3 -Instruct Mixtral Mixtral -Instruct Qwen2 Qwen2 -Instruct Qwen2 Qwen2 -Instruct
LP 79.6 77.1 82.5 82.3 80.4 79.7 85.4 85.4
MeanLP 77.1 74.8 79.6 79.4 77.7 77.1 74.5 74.3
PenLP 79.2 76.8 82.2 82.0 79.9 79.2 82.2 82.0
In-template LP $\mathbf{84.4}$±0.5 83.5±0.5 84.0±0.5 83.5±0.9 83.9±0.3 80.1±1.0 $\mathbf{87.9}$±0.3 $\mathbf{86.2}$±0.3
In-template MeanLP 82.6±0.7 81.9±0.5 82.6±0.3 82.2±0.8 82.0±0.7 78.7±1.1 77.7±1.2 77.5±1.4
In-template PenLP 83.8±0.5 83.0±0.5 83.8±0.4 83.3±1.0 83.2±0.4 79.8±1.1 83.4±0.4 82.9±0.5
In-template compar. LP 71.8±4.5 61.8±2.6 72.1±3.2 68.4±1.2 62.7±3.7 58.5±3.8 68.1±4.5 60.6±3.9
A/B prompting 77.4±3.6 81.9±3.7 76.5±4.3 80.5±3.5 80.8±1.1 82.5±0.3 77.3±4.2 80.9±1.6
Yes/No prob comp 73.6±3.2 $\mathbf{88.9}$±0.3 $\mathbf{84.1}$±1.2 $\mathbf{84.0}$±2.0 $\mathbf{89.0}$±0.2 $\mathbf{86.8}$±0.4 86.1±0.3 84.8±0.3
Table 3: Percentage accuracy (averaged over templates) by method and model. $\pm$ denotes standard deviation. The bold font denotes the best score. Underlines denote the second best. See Section C.1 for the max accuracy.

### v2

@@ Abstract
The grammatical knowledge of language models (LMs) is often measured using a benchmark of linguistic minimal pairs, where the LMs are presented with a pair of acceptable and unacceptable sentences and required to judge which is more acceptable. Conventional approaches directly compare sentence probabilities assigned by LMs, but recent large language models (LLMs) are trained to perform tasks via prompting, and thus, the raw probabilities they assign may not fully reflect their grammatical knowledge.
In this study, we attempt to derive more accurate acceptability judgments from LLMs using prompts and templates. Through extensive experiments in English and Chinese, we compare nine judgment methods and find two of them, a probability readout method—in-template LP and a prompt-based method—Yes/No probability computing, achieve higher accuracy than the conventional ones. Our analysis reveals that these methods excel in different linguistic phenomena, suggesting they access different aspects of LLMs’ knowledge. We also find that ensembling the two methods outperforms single methods. Consequently, we recommend these techniques, either individually or ensembled, as more effective alternatives to conventional approaches for assessing grammatical knowledge in LLMs. 111 Our codes and templates are published at https://github.com/Yusuke196/llm-acceptability.
@@
Yusuke Ide    Yuto Nishida    Justin Vasselli    Miyu Oba Yusuke Sakai    Hidetaka Kamigaito    Taro Watanabe Nara Institute of Science and Technology {ide.yusuke.ja6, nishida.yuto.nu8, vasselli.justin_ray.vk4, oba.miyu.ol2, sakai.yusuke.sr9, kamigaito.h, taro}@is.naist.jp
@@ 1 Introduction
The grammatical knowledge of language models (LMs) is often measured using acceptability judgments (Lau et al., 2017; Warstadt et al., 2019). There are two main categories of acceptability judgment benchmarks, the single-sentence one and the minimal-pair (MP) one, as detailed in Section 2. We focus on the latter because it allows us to directly measure LMs’ grammatical knowledge without task-specific fine-tuning. Below is an example of a minimal pair from Warstadt et al. (2020).
@@
Here, sentence (a) is acceptable or grammatically correct, while (b) is not, as its underlined verb violates the subject-verb agreement.
@@
Meanwhile, the recent scaling up of model sizes and training data for LMs has made it possible to solve a wide range of tasks using few-shot or zero-shot prompting, without the need for task-specific fine-tuning Brown et al. (2020); Liu et al. (2023), popularizing the term large language model (LLM). Incorporating learning techniques such as instruction tuning Wei et al. (2022) and Direct Preference Optimization (DPO) Rafailov et al. (2023) further improved the alignment of LLM outputs with human preferences. The LLMs trained by these techniques achieve good performance by prompting, i.e., guidance on what knowledge to elicit.
In this light, various methods can be developed to obtain more accurate acceptability judgments by providing LLMs with targeted guidance. However, as discussed in Section 2, most previous studies simply input the sentences into the (L)LM, calculate their probabilities, and consider the sentence with the higher probability as the acceptable one. Although Hu and Levy (2023) compared multiple methods of obtaining acceptability judgments from LLMs (see their Experiment 3b), their probability readout method and prompting method were limited to basic ones. As a result, they broadly claim that prompting is ineffective, whereas our experiments demonstrate that it can be highly effective with the proper technique.
@@
Figure 1: Conceptual illustration of our top methods. Differences between paired sentences are underlined. Both methods judge the sentence that results in a higher (log) probability acceptable. See Section 3 for the details.
We contribute to this area by comparing (1) the conventional sentence probability readout222 Readout refers to accessing an LLM’s output layer to compute probabilities of strings Kauf et al. (2024). methods, (2) our novel probability readout methods in in-template settings, and (3) prompt-based methods. In the in-template probability readout, we insert each sentence into a template and use the LLM to calculate the probability of the complete string. The template allows us to guide the LLM to judge the sentence’s grammaticality in a way that conventional probability readouts cannot. We call the most basic method in-template LP, where LP stands for log probability. For prompt-based methods, we investigate a basic method of asking LLMs to respond with a choice and Yes/No probability computing (Yes/No prob comp), where we compute the normalized probability of “Yes” versus “No”, inspired by UniEval Zhong et al. (2022). Figure 1 presents the conceptual illustration of in-template LP and Yes/No prob comp.
To rigorously compare these methods, we conduct experiments using eight LLMs and two MP benchmarks (one for English and one for Chinese). The results show the effectiveness of the two methods. In-template LP consistently outperforms conventional methods, achieving the highest accuracies on the Chinese benchmark. Yes/No prob comp achieves the highest accuracies on the English benchmark in all but one setting.
Moreover, our analysis demonstrates the following key findings. (1) In-template LP and Yes/No prob comp have different strengths; for example, Yes/No prob comp is robust against token-length bias. This indicates that they access different aspects of LLMs’ grammatical knowledge, contributing to a more comprehensive evaluation. (2) Ensembling the two methods further improves the accuracy, revealing their complementary capabilities. The highest score, achieved with Qwen2, is 1.6 percentage points higher than humans on the English benchmark. Based on these findings, we recommend the following: if possible, ensemble the two methods; otherwise, use in-template LP. (3) We identify a common weakness across all our combinations of the LLMs and methods: they struggle to make correct judgments on linguistic phenomena where the unacceptable sentence can be obtained by shuffling the words in the acceptable one, which presents a challenge for future work.
@@ 2 Related Work
Benchmarks of acceptability judgments can be divided into two categories: single-sentence benchmarks and MP benchmarks. The single-sentence benchmarks pose a binary classification of single sentences as seen in CoLA Warstadt et al. (2019), a dataset composed of sentences each labeled acceptable or unacceptable. CoLA was incorporated into the natural language understanding benchmark GLUE Wang et al. (2018) and has been used to evaluate a wide range of models, including LMs. However, single-sentence benchmarks cannot measure LMs’ grammatical knowledge directly because they require training a supervised classifier before the evaluation. This makes it difficult to distinguish between the knowledge of the model itself and what is learned by training the classifier Warstadt et al. (2020).
In contrast, MP benchmarks present minimally different pairs and the task is to determine which is the most acceptable sentence, eliminating the need for a classifier. As another advantage of the MP benchmark, minimal pairs can be automatically generated in a controlled manner, providing a sufficient amount of quality data for model evaluation Linzen et al. (2016). In conventional experiments using an MP benchmark, judgments are made based on sentence probabilities. Models are evaluated by whether they assign a higher probability to the acceptable sentence in each minimal pair. This method, which we call sentence probability readout, has been dominantly employed for MP acceptability judgments across languages (Marvin and Linzen, 2018; Warstadt et al., 2020; Mueller et al., 2020; Haga et al., 2024; Xiang et al., 2021; Someya and Oseki, 2023, inter alia).
Experiments with prompting LLMs have been conducted on both single-sentence benchmarks and MP benchmarks. Zhang et al. (2024) compared various models, including LLMs, on a single-sentence benchmark. On MP benchmarks, Hu and Levy (2023) compared the sentence probability readout and prompting. However, their probability readout and prompt-based methods relied on basic implementations, without systematically exploring alternative ways to calculate sentence probabilities or optimize prompting strategies.
Another line of work has studied biases that affect the performance of sentence probability readout. For example, it is known that the sentence probabilities given by LMs tend to decline as the token length of the sentence grows; normalized measures such as PenLP Wu et al. (2016) have been shown to mitigate some of this bias Lau et al. (2020), but they do not eliminate it Ueda et al. (2024).
@@
Table 1: Example English inputs of the readout methods. The target or inserted sentences are in italics. See Table 6 for Chinese versions.
@@
Table 2: Example English messages for prompting. The target or inserted sentences are in italics. See Table 7 for Chinese versions.
@@ 3.1 Sentence Probability Readout
We experiment with three measures to compute the probability scores: LP, MeanLP, and PenLP. All of them have been widely used in acceptability judgments.333 SLOR Kann et al. (2018) is also commonly used, but it requires building a unigram model using the training data of the LM. Because the training data of the LLMs we use is not publicly available, we skip the acceptability measure. LP is the unnormalized log probability of the sentence
$\mathrm{LP}(\bm{s})$ $=\log P(\bm{s})$ (1)
where $\bm{s}$ is the input sequence of tokens and $P(\bm{s})$ is the probability assigned to $\bm{s}$ by the model
$P(\bm{s})$ $=\prod_{i=1}^{|\bm{s}|}P(t_{i}|t_{<i}).$ (2)
@@
where $\alpha$ is a hyperparameter to scale the sentence length, reducing the impact of long sentences and ensuring a fair comparison across different lengths. We set $\alpha=0.8$ following Lau et al. (2020) and Ueda et al. (2024). We hereafter refer to the three judgment methods simply by the name of the corresponding measures: LP, MeanLP, PenLP.
@@ 3.2 In-template Probability Readout
In-template probability readout follows the same steps of computing and comparing probabilities as sentence probability readout. Meanwhile, its input string is built by embedding the sentences in a template designed to draw focus to their grammaticality. The input has two types: in-template single and in-template comparative. For each type, we prepare five templates per language because the performance can vary due to minor differences in expressions within prompts Zheng et al. (2023). The templates were created based on those of Flan444 https://github.com/google-research/FLAN Wei et al. (2022). For Chinese experiments, we use translations of English templates. Translations were generated by DeepL [^5] https://www.deepl.com and post-edited by a native Chinese speaker.
@@
As the length of the input changes depending on the length of the target sentence, predictions by in-template single inputs must also use normalization techniques. We thus apply each of the three measures explained above to the method, dubbing the corresponding methods in-template LP, in-template MeanLP, and in-template PenLP, respectively. The final measure depends on whether we base our computations on the whole input string or the target sentence only. We report the result of the former because it performed better in our preliminary experiments.
@@
In-template comparative inputs are built by filling two placeholders; we insert the target sentence into the first one and the other sentence of the minimal pair into the second. Table 1 shows an example input. Note that the second sentence is supplementary, and our aim here is to measure the acceptability of the first one.
In-template comparative does not need normalization, because the sum of the token lengths of two sentences and, thus, the length of the whole input string is constant, no matter which of the paired sentences comes first. Hence, we only calculate LP for the in-template comparative input, referring to this method as in-template comparative LP.
3.3 Prompt-based Methods
@@
A/B prompting inputs a prompt containing the paired sentences to the models and asks which sentence is acceptable. The prompt is exemplified in Table 2. The user message contains one acceptable and one unacceptable sentence. Their order (which sentence goes to A or B) is randomized to eliminate the potential bias from the order Pezeshkpour and Hruschka (2023). We perform constrained decoding by outlines777 https://github.com/outlines-dev/outlines Willard and Louf (2023) to ensure that the model outputs either A or B, because our preliminary experiments without outlines observed many outputs violating the constraint. We turn off sampling in decoding.
@@
In Yes/No prob comp, we compute the score of each sentence as the normalized probability of “Yes” versus “No” given a prompt asking its acceptability. An example prompt is shown in Table 2. We predict the sentence that resulted in a higher “Yes” probability to be acceptable. This method is inspired by UniEval Zhong et al. (2022), which is shown to correlate well with human judgments in evaluating natural language generation. We formulate the probability given a sentence $\bm{s}$ as follows,
@@
where $P_{\mathrm{LLM}}(\cdot)$ is the probability of a word assigned by the model. For the Chinese experiments, we substitute “是” and “否” for “Yes” and “No”, respectively, if these words are not segmented into subwords; otherwise, we employ the same formulation as English experiments.
Llama-3 Llama-3-Inst. Mixtral Mixtral-Inst. Qwen2 Qwen2-Inst.
LP 79.6 77.1 82.5 82.3 80.4 79.7
MeanLP 77.1 74.8 79.6 79.4 77.7 77.1
PenLP 79.2 76.8 82.2 82.0 79.9 79.2
In-template LP $\mathbf{84.4}$±0.5 83.5±0.5 84.0±0.5 83.5±0.9 83.9±0.3 80.1±1.0
In-template MeanLP 82.6±0.7 81.9±0.5 82.6±0.3 82.2±0.8 82.0±0.7 78.7±1.1
In-template PenLP 83.8±0.5 83.0±0.5 83.8±0.4 83.3±1.0 83.2±0.4 79.8±1.1
In-template compar. LP 71.8±4.5 61.8±2.6 72.1±3.2 68.4±1.2 62.7±3.7 58.5±3.8
A/B prompting 77.4±3.6 81.9±3.7 76.5±4.3 80.5±3.5 80.8±1.1 82.5±0.3
Yes/No prob comp 73.6±3.2 $\mathbf{88.9}$±0.3 $\mathbf{84.1}$±1.2 $\mathbf{84.0}$±2.0 $\mathbf{89.0}$±0.2 $\mathbf{86.8}$±0.4
(a) BLiMP
Llama-3 Llama-3-Inst. Yi-1.5 Yi-1.5-Chat Qwen2 Qwen2-Inst.
LP 83.2 80.4 86.8 85.3 85.4 85.4
MeanLP 74.5 71.7 75.9 74.5 74.5 74.3
PenLP 80.3 77.7 84.3 82.3 82.2 82.0
In-template LP $\mathbf{85.7}$±0.5 $\mathbf{82.9}$±0.3 $\mathbf{87.4}$±0.4 $\mathbf{86.8}$±0.8 $\mathbf{87.9}$±0.3 $\mathbf{86.2}$±0.3
In-template MeanLP 79.9±0.9 77.7±0.6 78.8±1.3 79.4±1.2 77.7±1.2 77.5±1.4
In-template PenLP 83.0±0.5 80.7±0.4 84.0±0.8 83.3±1.1 83.4±0.4 82.9±0.5
In-template compar. LP 68.2±2.4 58.2±3.0 63.9±4.6 61.8±3.4 68.1±4.5 60.6±3.9
A/B prompting 68.0±3.5 69.2±4.0 74.2±3.6 75.5±3.0 77.3±4.2 80.9±1.6
Yes/No prob comp 76.3±1.3 76.9±0.9 78.2±0.9 81.7±0.3 87.2±0.4 83.9±0.4
(b) CLiMP
Table 3: Percentage accuracy (averaged over templates) by method and model. $\pm$ denotes standard deviation. The bold font denotes the best scores. Underlines denote the second best. See Section D.1 for the max accuracy. * denotes scores significantly higher than LP ($p\leq 0.01$).